# Knapsack Optimization-based Schema Linking for LLM-based Text-to-SQL Generation

### v1

@@ 6 Related Work
Schema linking is essential for aligning natural language queries with database components, playing a pivotal role in the overall text-to-SQL process. Recent approaches utilize large language models (LLMs) to enhance schema linkage through methods such as semantic retrieval, historical query integration, and bidirectional linking (Pourreza and Rafiei, 2024; Cao et al., 2024; Zhang et al., 2024). Techniques like DAIL-SQL and E-SQL refine linkage by leveraging LLMs’ contextual understanding and query enrichment capabilities (Gao et al., 2024; Caferoğlu and Ulusoy, 2024). Retrieval-based strategies focus on retrieve information from database to improve SQL generation (Pourreza and Rafiei, 2024). Chess (Talaei et al., 2024) and CodeS (Li et al., 2024b) utilize semantic retrieval to improve schema linking by accessing relevant database information. Despite these advancements, many state-of-the-art text-to-SQL models emphasize fine-tuning strategy. DTS-SQL (Pourreza and Rafiei, 2024) fine-tunes a local LLM to select relevant tables from user queries, while Solid-SQL (Liu et al., 2025) introduces a data augmentation method for enhancing schema linking. DELLM (Hong et al., 2024a) finetuned a specialized data expert LLM to provide additional knowledge for schema linking and SQL generation. Approaches like TA-SQL (Qu et al., 2024) use in-context learning to guide LLMs, generating dummy SQL for linked element abstraction. Similarly, RSL-SQL (Cao et al., 2024) employs a bidirectional approach for schema simplification. Meanwhile, E-SQL (Qu et al., 2024) and (Maamari et al., 2024) note that schema linking challenges persist, leading some models to use the entire schema directly as input for SQL generation.
@@ 7 Conclusion
This paper introduced the Knapsack Schema Linking Agent (KaSLA), a pioneering approach to overcoming schema linking challenges in text-to-SQL tasks. Based on knapsack optimization, KaSLA effectively prevent the missing of relevant elements and excess redundancy, thereby significantly enhancing SQL generation accuracy. Our introduction of an enhanced schema linking metric, sets a new standard for evaluating schema linking performance. KaSLA utilizes a hierarchical linking strategy, starting with optimal table linking and then proceeding to column linking within selected tables, effectively reducing the candidate search space. In each step, KaSLA employs a knapsack optimization strategy to link potentially relevant elements while considering a limited tolerance for potential redundancy. Extensive experiments on the Spider and BIRD benchmarks have demonstrated that KaSLA can significantly enhance the SQL generation performance of existing text-to-SQL models by substituting their current schema linking processes. These findings underscore KaSLA’s potential to revolutionize schema linking and advance the broader capabilities of text-to-SQL systems, offering a robust solution that paves the way for more precise and efficient database interactions.
8 Limitation
There are mainly two limitations of this work. First, although KaSLA demonstrates significant advancements in schema linking accuracy and efficiency, its performance traditionally depends on a comprehensive training dataset with detailed ground truth linking results. However, our experiments indicate that KaSLA possesses a degree of transferability, showing promise even in scenarios where training data is sparse or less diverse. This adaptability suggests that KaSLA can maintain reasonable effectiveness across different database environments, although performance might still be impacted in highly dynamic or rapidly evolving schema structures. Additionally, the reliance on a fine-tuning process of LLMs, integrating DeepSeek-coder-1.3B, could pose scalability challenges in resource-constrained settings. Despite these limitations, KaSLA remains a powerful tool for schema linking optimization in text-to-SQL tasks, and future work could focus on enhancing its adaptability and efficiency.
9 Ethics Statement
We confirm that we have fully complied with the ACL Ethics Policy in this study. All the datasets are publicly available and have been extensively used in research related to text-to-SQL.
Appendix A Implementation Details.
We implemented the proposed KaSLA and baselines under the following settings: For top-$K$ used in the redundancy tolerance estimation of KaSLA, we set it to 30. For method using fine-tuned LLM, We utilized LoRA (Hu et al., 2021) for parameter-efficient fine-tuning. The learning rate was initialized to $41e-$ 4 with a cosine decay, the batch size was set to 16, and the training epoch was 3.
Appendix B Input formats
We provided the input and output formats of the binary scoring function of KaSLA introduced in section 4.1 in Table 6.
Type Input format
Full schema table department, columns = [ department__\__id (int |||| primary key||||values: 1), name (text||||values: State, Treasury), ranking (int |||| values: 1, 2)] table head, columns = [ head__\__id (int |||| primary key |||| values: 1), name (text |||| values: Tiger Woods), age (real |||| values: 67.0)] table management, columns = [ department__\__id ( int |||| primary key |||| values: 2), head__\__id (int |||| primary key |||| values: 5), temporary__\__acting (text |||| values: Yes)] foreign keys: management.head__\__id = head.head__\__id, management.department__\__id = department.department__\__id
Question How many heads of the departments are older than 56? Generate SQL to solve the above question and list the relevant columns in each table
Type Output format
Simulated SQL SELECT count(*) FROM head WHERE age > 56
Schema linking results "department": null, "management": null, "head": [ "age" ]
Table 6: Input and output formats of the binary scoring model.
Appendix C The Dynamic Programming Approach used in KaSLA
We illustrate the details of the dynamic programming algorithm used in KaSLA in Algorithm 1.
Input: relevance score set ${\mathcal{R}}$, redundancy score set ${\mathcal{W}}$, the upper redundancy tolerance ${\mathcal{U}}$
Output: selection set $\hat{\mathcal{H}}$
$n\leftarrow|{\mathcal{R}}|$, $^{}^\hat{\mathcal{H}}\leftarrow\{\}$
[…]
$keep\leftarrow\text{array of }(n+1)\times({\mathcal{U}}+1)$
initialized to False
for $i\leftarrow 1$ to $n$ do
       for $w\leftarrow 0$ to ${\mathcal{U}}$ do
             if ${\mathcal{W}}[i-1]\leq w$ then
                         $keep[i][w]\leftarrow\text{True}$
                  else
                         $\mathcal{A}[i][w]\leftarrow\mathcal{A}[i-1][w]$
            else
                   $\mathcal{A}[i][w]\leftarrow\mathcal{A}[i-1][w]$
$k\leftarrow{\mathcal{U}}$  for $i\leftarrow n$ downto $1$ do
       if $keep[i][k]$ then
             ^.$\hat{\mathcal{H}}.\text{add}(i-1)$
             $k\leftarrow k-{\mathcal{W}}[i-1]$
return $\hat{\mathcal{H}}$
Algorithm 1 Dynamic Programming-based 0-1 Knapsack Problem Optimization
Appendix D Complexity Analysis
The efficiency of KaSLA stems from its hierarchical approach, dynamic programming and inherent parallelism. KaSLA’s time complexity is determined by two main phases: table linking and column linking within the linked tables. For a database with $n_{t}$ tables and a table capacity of $C_{t}$, the table linking phase has a time complexity of $O(n_{t}C_{t})$. KaSLA then links columns within the selected table. Let $n_{c}$ and $C_{c}$ represent the number of columns and the column capacity of a selected table, respectively. The time complexity of the column linking phase is $O(n_{c}C_{c})$. Consequently, the total time complexity of KaSLA is $O(n_{t}C_{t}+n_{t}(n_{c}C_{c}))$, indicating that the complexity of KaSLA scales relevant with the database dimensions.
Appendix E The Ablation Studies for Different Language Models
We present the ablation studies on the choice of language models for the probabilistic scoring model, as well as the performance scaling of various language models for the binary scoring model used in KaSLA. These studies are detailed in Table 8 and Table 7, respectively.
Dataset BERT -Base BERT -Large RoBERTa -Base RoBERTa -Large SGPT -1.3B
BIRD-dev 59.12 59.52 59.65 60.17 59.97
Table 7: Execution Accuracy (EX) (%) of the scaling up performance about the choice of language model for binary scoring model. We conduct the experiments for CodeS + KaSLA on BIRD-dev and use CodeS with CodeS-15B as the text-to-SQL backbone.
Dataset Deepseek-coder -1.3B Deepseek-coder -6.5B Deepseek-coder -7B
BIRD-dev 60.17 61.28 61.47
Table 8: Execution Accuracy (EX) (%) of the ablation study about the choice of language model for probabilistic scoring model. We conduct the experiments for CodeS + KaSLA on BIRD-dev and use CodeS with CodeS-15B as the text-to-SQL backbone.

### v2

@@ VI-B Schema Linking in Text-to-SQL
Within the text-to-SQL task, schema linking is proven as an essential module for aligning natural language queries with database contents, playing a pivotal role in the overall SQL generation process [3]. Recent approaches utilize LLMs to enhance schema linkage through methods such as semantic retrieval, historical query integration, and bidirectional linking [9, 12, 41]. Techniques like DAIL-SQL [4] and E-SQL [26] refine linkage by leveraging LLMs’ contextual understanding and query enrichment capabilities [4, 26]. Retrieval-based strategies focus on retrieving information from the database to improve SQL generation [9]. CodeS [8] utilizes semantic retrieval to improve schema linking by accessing relevant database information. CHESS [11] design dedicated schema selector agents that prompt LLMs to choose relevant tables and columns. Despite these advancements, many state-of-the-art text-to-SQL models emphasize fine-tuning strategy. DTS-SQL [9] fine-tunes a local LLM to select relevant tables from user queries, while Solid-SQL [10] introduces a data augmentation method for enhancing schema linking. Approaches like TA-SQL [5] use in-context learning to guide LLMs, generating dummy SQL for linked element abstraction. Similarly, RSL-SQL [12] employs a bidirectional approach for schema simplification. Meanwhile, E-SQL [5] and Distillery-SQL [6] note that schema linking challenges persist, leading some models to use the entire schema directly as input for SQL generation.
@@ VII Conclusion
This paper studies the schema linking in text-to-SQL, specifically how to link relevant elements while avoiding missing and redundant ones. We first proposed Recall+ and Precision+ with a restricted missing-element indicator to better assess linking performance. Furthermore, we introduced the Knapsack Schema Linking Approach (KaSLA), a pioneering approach to overcoming schema linking challenges in text-to-SQL tasks. Based on knapsack optimization, KaSLA effectively links relevant elements while preventing missing and redundant elements simultaneously, thereby significantly improving SQL generation accuracy. KaSLA utilizes a hierarchical linking strategy, starting with optimal table linking and then proceeding to column linking within selected tables, effectively reducing the candidate search space. In each step, KaSLA employs a knapsack optimization strategy to link potentially relevant elements while considering a limited tolerance for potential redundancy. Extensive experiments on the Spider and BIRD benchmarks have demonstrated that KaSLA can significantly enhance the SQL generation performance of existing text-to-SQL models by substituting their current schema linking processes. These findings underscore KaSLA’s potential to revolutionize schema linking and advance the broader capabilities of text-to-SQL systems, offering a robust solution that paves the way for more precise and efficient database interactions.